APPENDIX D: Some Sample Pages From a Survey Typeset in TEX This appendix consists of TEX output for several selected pages of a survey recently conducted by a University of Oregon graduate student. It is included as an example of how to use TEX to typeset typical survey instruments. A machine-readable copy of the raw TEX commands used to typeset these pages is available to OREGON VAX users as:

      TEX$DISK:[TEX.SAMPLES]SURVEY.TEX

On other systems get a copy of that file via anonymous FTP.

FTP to DECOY.UOREGON.EDU (128.223.32.19) and see directory pub/tex/samples.

THE \underline{TOURISM} INDUSTRY

6. TOLERABLE REDUCTION IN TOURISM INDUSTRY \underline{JOBS}: The tourism and recreation industries provide about 21,500 direct jobs and 20,300 indirect jobs in Oregon.

Assume that in order to maintain current timber harvest levels, a certain percentage of those tourism and recreation jobs may get eliminated.

What percentage of tourism industry jobs would you be willing to eliminate (if necessary) in order to maintain the current timber harvest levels?

0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%

MAXIMUM TOLERABLE REDUCTION IN TOURISM INDUSTRY JOBS

7. TOLERABLE TOURISM INDUSTRY \underline{HOME FORECLOSURES}: Elimination of tourism and recreation jobs in Oregon would probably have a number of economic consequences, including increasing the number of tourism and recreation industry employee home mortgages which fail.

What percentage of tourism industry home foreclosures would you be willing to accept (if necessary) in order to maintain the current timber harvest levels?

0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%

MAXIMUM TOLERABLE TOURISM INDUSTRY HOME FORECLOSURES

8. SPENDING FOR DISPLACED TOURISM WORKER \underline{SUPPORT}: Assume that some people formerly employed by the tourism industry would be unable to find new work, if tourism and recreation jobs are reduced.

It has been proposed that tax money could be used to help support those displaced tourism and recreation workers until they could find new work.

How much extra would you be willing to pay next year (in Oregon income tax) to directly support tourism and recreation industry workers who can’t find new jobs?

$0 $25 $50 $75 $100 $125 $150 $175 $200 OTHER: $

9. SPENDING FOR DISPLACED TOURISM WORKER \underline{RETRAINING}: Suppose the people who might lose their tourism and recreation industry jobs needed retraining in order to be able to find new work. Assume that this sort of retraining expense would be paid for via an increase in Oregon’s income tax.

How much extra would you be willing to pay next year (in Oregon income tax) to support retraining tourism workers who can’t find new jobs?

$0 $25 $50 $75 $100 $125 $150 $175 $200 OTHER: $

THE ENVIRONMENT vs. \underline{TIMBER} JOBS

11(a). What’s the maximum reduction in timber industry jobs which you’d be willing to accept in order to maintain Oregon’s current level of \underline{clean water}? 0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%

MAXIMUM TOLERABLE REDUCTION IN TIMBER INDUSTRY JOBS

11(b). What’s the maximum reduction in timber industry jobs which you’d be willing to accept in order to maintain Oregon’s current level of \underline{flood control} (flood control holds water in the forest and prevents run-off that could flood inhabited areas)? 0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%

MAXIMUM TOLERABLE REDUCTION IN TIMBER INDUSTRY JOBS

11(c). What’s the maximum reduction in timber industry jobs which you’d be willing to accept in order to maintain Oregon’s current level of \underline{clean air}? 0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%

MAXIMUM TOLERABLE REDUCTION IN TIMBER INDUSTRY JOBS

11(d). What’s the maximum reduction in timber industry jobs which you’d be willing to accept in order to maintain Oregon’s current level of \underline{scenic beauty along highways}? 0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%

MAXIMUM TOLERABLE REDUCTION IN TIMBER INDUSTRY JOBS

11(e). What’s the maximum reduction in timber industry jobs which you’d be willing to accept in order to maintain Oregon’s current level of \underline{wilderness areas}? 0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%

MAXIMUM TOLERABLE REDUCTION IN TIMBER INDUSTRY JOBS

11(f). What’s the maximum reduction in timber industry jobs which you’d be willing to accept in order to maintain Oregon’s current level of \underline{erosion control} (holding top soil in place and preventing mud slides and the muddying of streams and rivers)? 0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%

MAXIMUM TOLERABLE REDUCTION IN TIMBER INDUSTRY JOBS

11(g). What’s the maximum reduction in timber industry jobs which you’d be willing to accept in order to maintain Oregon’s current level of \underline{wild fish runs}? 0% 10% 20% 30% 40% 50% 60% 70% 80% 90% 100%

MAXIMUM TOLERABLE REDUCTION IN TIMBER INDUSTRY JOBS

MAGAZINES

24. What magazines do you normally read? (check all that apply) 1 General News Magazines (Time, Newsweek, etc.) 2 Sports Magazines (Sports Illustrated, Golf, Scuba, etc.) 3 Outdoors Magazines (Field and Stream, Guns and Ammo, etc.) 4 Farm/Forestry Magazines (Farm Journal, Successful Farming, etc.) 5 Environmental Magazines (Audubon, Sierra, Conservationist, etc.) 6 Technical/Professional Magazines (NE J. of Medicine, Proc. IEEE, etc.) 7 General Science Magazines (Scientific American, Nature, etc.) 8 General Interest Magazines (Life, Reader’s Digest, etc.) 9 Women’s Magazines (McCalls, House Beautiful, Cosmopolitan, etc.) A Men’s Magazines (Playboy, Penthouse, etc.) B Natural History Magazines (National Geographic, Smithsonian, etc.) C Literary Magazines (New Yorker, Harper’s, etc.) D Supermarket tabloid-format papers (National Enquirer, etc.)

NEWSPAPERS

25. Do you regularly get the \underline{LOCAL} paper? 1 YES, I have the local paper delivered to me. 2 YES, I usually get the local paper from a newspaper box, or store. 3 SOMETIMES I’ll pick up the local paper, but usually I don’t. 4 NO, I rarely (or never) get the local paper.

26. Do you regularly get a \underline{NATIONAL} paper (such as the N.Y. Times?) 1 YES, I have a national paper delivered to me. 2 YES, I usually get a national paper from a newspaper box or store. 3 SOMETIMES I’ll pick up a national paper, but usually I don’t. 4 NO, I rarely (or never) get a national newspaper.